OpenStudy (anonymous):

A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle Θ. The area A of the opening may be expressed as the function: A(Θ) = 16 sin Θ (cos Θ + 1). If Θ = 60°, what is the area of the opening?

OpenStudy (anonymous):

A(45°) = 16 sin(45°)(cos(45°) + 1)    = 16(1/√2)(1 + 1/√2)    = 16(1/√2)(1 + √2)/√2    = 16(1 + √2)/2    = 8 + (√2)/2 in²

OpenStudy (anonymous):

Can you answer more question?

OpenStudy (anonymous):

ya make another question and tag me